OpenStudy (anonymous):

College Algebra: Functions and Graphs: Find x- and y-intercepts: -1/2x+y-4=0 Stuck on finding y-intercept, I KNOW HOW TO SOLVE, I MUST OF JUST MADE A SIMPLE MISTAKE... My work so far: AX+BY+C=0 X-int; y=0 -1/2X+0-4=0 -1/2x=4 x=-2, x-int: (-2,0) *** y-int; x=0 -1/2(0)+y-4=0 0+y-4=0 y=4; y-int: (0,4) <--- answer key says its (-8,0) though... how??

OpenStudy (amistre64):

-1/2x + y - 4 = 0 the y intercepts is a point on the y axis (also called the x=0 axis) so let x=0 y - 4 = 0, and y=4 your process was fine, just make sure you are matching the correct problem in the answer key maybe?

OpenStudy (anonymous):

@amistre64 so was my answer for the y-int correct? (0,4)?

OpenStudy (amistre64):

given: -1/2 x+y-4=0 ^^ x not a denominator im assuming then yes, the y intercept is the point (x=0, y=4)

OpenStudy (amistre64):

given: -1/2 x+y-4=0 the x intercecpt is the point (x=-8,y=0)

OpenStudy (anonymous):

yes not a denominator okay thank you!! :)

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

@amistre64 I keep getting (-2,0) for the x-int though. my work: -1/2x+y-4=0 -1/2x+0-4=0 -1/2x-4=0 -1/2x=4 -1/2x-/1/2=4/-1/2 x=-2

OpenStudy (amistre64):

does --2/2 = 4?

OpenStudy (amistre64):

-1/2x=4, multiply by -2 --2/2 x=4(-2) x = 4(-2)

OpenStudy (amistre64):

i think you were mentally messing yourself up with that. you knew -2 was a part of the process and just skipped the final step

OpenStudy (amistre64):

or you forgot how to divide by a fraction

OpenStudy (amistre64):

4 ---- = 4(-2/1) -1/2

OpenStudy (anonymous):

yep def forgot to divide by a fraction... I see it now. thank you so much for making it clear to me!

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

Thanks a billion!!!!